OpenStudy (anonymous):

Determine the infinite limit? (I have to present it in class. can I have explanation please). lim + x + 2 x->-3 ------- = x + 3

OpenStudy (tkhunny):

You might try long division. \(\dfrac{x+2}{x+3} = 1 - \dfrac{1}{x+3}\) The numerator is a whole lot less confusing in this form.

OpenStudy (anonymous):

Or somewhat differently, \[\frac{x+2}{x+3}=\frac{x+2}{x+3}\cdot\frac{\frac{1}{x}}{\frac{1}{x}}=\frac{1+\frac{2}{x}}{1+\frac{3}{x}}\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."